### v3

# Spatiotemporal Photon Blockade for Nonreciprocal Quantum Absorption

Abstract
Controlling the flow of photons is crucial for advancing quantum technologies. We introduce the concept of spatiotemporal photon blockade for nonreciprocal quantum absorption, utilizing space-time-periodic metasurfaces. Our study presents a methodology for experimentally realizing this effect, where photon frequency coherence with the metasurface’s space-time modulation enables one-way quantum absorption. In this system, forward-traveling photons are energetically modulated and absorbed within the slab, while backward-traveling photons are transmitted without interaction. Our analysis includes band structure, isofrequency diagrams, and nonreciprocal absorption results. These findings lay the groundwork for developing nonreciprocal quantum devices and enhancing photon management in milli-Kelvin temperature quantum systems.
Photon blockade is a quantum optical phenomenon analogous to the Coulomb blockade, where the suppression of electrical transport occurs due to electronic repulsion, preventing multiple electrons from occupying the same region. In photon blockade, the absorption or transmission of a single photon inhibits the absorption or transmission of subsequent photons. This phenomenon occurs in systems where strong nonlinear interactions between photons effectively create a scenario in which only one photon can be absorbed or transmitted at a time, leading to quantized light behavior. This effect, typically observed in systems such as atoms, quantum dots, or superconducting circuits coupled to resonant cavities, involves the system entering an excited state after the first photon interaction. This state shift prevents further photon interactions until the system returns to its ground state, which is crucial for quantum information processing and communication. Historically, photon blockade has been achieved in systems with strong coupling between quantum emitters and optical cavities, where the nonlinearity creates an energy gap sufficient to block additional photons  (?, ?). Recent advancements have expanded this concept to nonreciprocal photon blockade in rotating nonlinear devices, where device geometry or external fields control the directionality of photon interactions  (?). Additionally, loss-induced nonreciprocal photon blockade has been demonstrated, where engineered losses enable nonreciprocal photon dynamics, enhancing control over photon flow in quantum circuits  (?, ?).
Quantum nonreciprocity is increasingly relevant for superconducting circuits, which are essential for scalable quantum computing and communication networks. Nonreciprocal elements like isolators and circulators are critical for protecting qubits from back-reflected signals and routing quantum information. Recent studies have explored various mechanisms for achieving quantum nonreciprocity, including dynamic modulation and parametric interactions in superconducting devices  (?, ?). Observations of superconducting nonreciprocity through diamagnetic effects and engineered circuit asymmetries offer robust platforms for nonreciprocal quantum devices  (?, ?, ?, ?). Traditional directional absorbers and electronic components, such as varactors, transistors, and diodes, face limitations and introduce noise at millikelvin temperatures typical of quantum technologies. To overcome these challenges, we propose a novel approach for nonreciprocal photon blockade through spatiotemporal modulation in quantum metasurfaces. By leveraging the unique properties of superconducting materials and dynamically modulating the metasurface structure, we achieve directional quantum absorption, where photons are selectively absorbed or transmitted based on their direction of incidence. This spatiotemporal photon blockade presents new opportunities for controlling quantum light-matter interactions and offers significant advancements for quantum information technologies and nonreciprocal photonic devices. Our approach achieves efficient directional absorption with simultaneous amplification, a rare combination in conventional devices. It addresses inefficiencies and spurious signal generation, paving the way for high-performance signal processing applications.
Space-time refractive index modulation is conventionally realized through the modulation of electrical permittivity, as it is challenging to spatio-temporally modulate the magnetic permeability of materials  (?). In  (?, ?), it was shown for the first time that modulating both permeability and permittivity results in physical phenomena that cannot be achieved through conventional permittivity-modulated media, including zero reflection, zero band-gap, a large nonreciprocal response, and strong space-time harmonics. Furthermore, our metasurface demonstrates nonreciprocal behavior, with different absorption responses for incident waves from opposing directions, emphasizing the unidirectional properties imparted by the spatiotemporal modulation  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and time-varying systems (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). Utilizing analytical techniques such as Bloch-Floquet solutions, we provide a comprehensive understanding of the wave scattering phenomena from our metasurface, validating its design and offering insights for further optimization. Unlike linear systems constrained by phase-matching and dispersion, our nonlinear spatiotemporal media achieve superior performance in a compact form, making them ideal for quantum technologies.
1 Spatiotemporal Photon Blockade
Consider photons with frequency $\omega_{0}$ incident on a slab with a space-time periodic current density $J(z,t)$, modulated at a temporal frequency $\omega_{\text{s}}$, where $\omega_{\text{s}}=\omega_{0}$. We demonstrate that the temporal coherence between the incident photons and the periodic modulation of the superconducting slab induces a nonreciprocal photon blockade. This phenomenon arises because the energy associated with the space-time modulation effectively inhibits the transmission of subsequent photons. Instead, the incident photons transition to higher energy states within the slab and are absorbed, preventing their exit. This results in a one-way quantum absorption process, facilitated by the coherent spatiotemporal modulation of the superconducting slab.
Figure 1: Nonreciprocal photon blockade in superconducting spatiotemporal metasurfaces, where the space-time modulation and incident photons share the same frequency ($\omega_{\text{s}}=\omega_{0}$). The system comprises space-time-modulated nonlinear unit cells coupled together. a) Incidence from the left port ($L_{\text{in}}$): Photon blockade and energy transition to higher states occur due to the strong interaction and one-way coherency with the nonlinear left-to-right traveling space-time modulation, resulting in zero transmission to the right port ($R_{\text{out}}=0$). b) Illustration of the physical mechanism for incidence from the right port ($R_{\text{in}}$): Free passage of photons to the left port ($L_{\text{out}}=R_{\text{in}}$) occurs due to the lack of transition to higher energy states and weak interaction with the opposite-direction traveling nonlinear space-time modulation, resulting in full transmission of photons.
Figure 1 illustrates the principle of spatiotemporal photon blockade enabling nonreciprocal quantum absorption by leveraging the interplay between the incident photons and the one-way coherent dynamic modulation. Figure 1a demonstrates the photon blockade effect when photons are incident from the left port ($L_{\text{in}}$). Here, the photons interact strongly with the space-time modulation traveling from left to right. This interaction causes the system to enter an excited state, leading to energy transitions and blocking the transmission of photons to the right port ($R_{\text{out}}=0$). In contrast, Fig. 1b illustrates the scenario for photons incident from the right port ($R_{\text{in}}$). In this case, the interaction with the space-time modulation traveling in the opposite direction is weak. Consequently, there is minimal energy transition and the photons are transmitted freely to the left port ($L_{\text{out}}=R_{\text{in}}$). This nonreciprocal behavior, where the photon transmission and absorption are direction-dependent, underscores the potential of spatiotemporal metasurfaces for applications in quantum communication and sensing.
Figure 2: One-way photon blockade and absorption functionality.
Figure 2 illustrates the proposed one-way photon blockade and absorption mechanism facilitated by the coherently driven quantum spatiotemporal metasurface and its functionality. The metasurface operates as a nonreciprocal element, permitting photon transmission in one direction while blocking it in the opposite direction. To understand the nonreciprocal behavior of the spatiotemporal quantum metasurface, we first analyze its band structure using the dispersion relation. The analysis reveals crucial insights into the energy transitions and nonreciprocity induced by the space-time modulation. Figure 3a presents the $\omega-\kappa$ diagram, showcasing the band structure of the metasurface at the critical point where $\omega_{0}/\omega_{\text{s}}=1$. This diagram illustrates a strong nonreciprocity in the system. Specifically, the higher-order harmonics that travel in the forward $+z$ direction converge at the normalized wave vector $\kappa_{n}/\kappa_{\text{s}}=1$. This convergence indicates a resonant interaction between the incident photons and the space-time-modulated medium, leading to the absorption of energy from the fundamental harmonic. The absorbed energy causes a transition from the fundamental frequency $\omega_{0}$ to higher energy states denoted as $\omega_{m}$.
Figure 3: Band structure of the spatiotemporal quantum metasurface computed. (a) The $\omega-\kappa$ diagram illustrates strong nonreciprocity at $\omega_{0}/\omega_{\text{s}}=1$. Here, all higher-order forward ($+z$ direction) traveling space-time harmonics converge at $\kappa_{n}/\kappa_{\text{s}}=1$, resulting in the absorption of energy from the fundamental harmonic and transitioning the system from $\omega_{0}$ to higher frequency states, denoted as $\omega_{m}$. In contrast, the higher-order backward ($-z$ direction) traveling space-time harmonics are spatially separated, preventing any energy transition from $\omega_{0}$ to higher states. (b) The $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$ further demonstrates the nonreciprocity of the metasurface, where all forward higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$, effectively absorbing the energy of the fundamental harmonic. The group velocity vector $v_{\text{g},n}$ indicates that these harmonics propagate along the $+z$ direction, parallel to the metasurface boundary, rather than being transmitted outside the metasurface.
In contrast, the higher-order harmonics that travel in the backward $-z$ direction are spatially separated in the $\omega-\kappa$ diagram. This separation prevents any significant interaction between the incident photons and the medium, thereby inhibiting the energy transition from $\omega_{0}$ to higher states. As a result, the backward-propagating waves do not experience the same degree of energy absorption as the forward-propagating waves, which directly contributes to the observed nonreciprocal behavior of the metasurface. Figure 3b shows the $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$. This diagram further confirms the nonreciprocal nature of the metasurface, where all forward-propagating higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$. The clustering of these harmonics facilitates the absorption of the fundamental harmonic’s energy, while the group velocity vector $v_{\text{g},n}$ indicates that the energy is predominantly confined and directed along the $+z$ axis, parallel to the metasurface boundary. This confinement prevents the transmission of energy to the exterior of the metasurface, leading to strong nonreciprocal absorption.
The phenomena observed in the band structure is related to the concept of spatiotemporal photon blockade. In photon blockade, the presence of a single photon can prevent the passage of subsequent photons due to nonlinear interactions, effectively ”blocking” further photon transmission. In our system, this blockade effect is realized through the spatiotemporal modulation of the metasurface. The space-time modulation, with a frequency $\omega_{\text{s}}$ matched to the incident photon frequency $\omega_{0}$, induces a strong nonlinear interaction for forward-propagating waves. As a result, when photons from the left are incident on the metasurface, they are resonantly coupled to the modulation, leading to energy absorption and transition to higher frequency states (spatial harmonics) $\omega_{m}$. This absorption effectively prevents the transmission of these photons through the metasurface, analogous to a photon blockade. The nonreciprocity arises because the same modulation does not facilitate similar energy transitions for backward-propagating waves, allowing them to pass through the metasurface without significant energy loss. Thus, the metasurface behaves as a nonreciprocal photon blockade device, where the space-time modulation selectively blocks photons traveling in one direction by shifting their energy to higher states, while allowing photons traveling in the opposite direction to pass freely. This behavior is crucial for potential applications in quantum information processing and nonreciprocal quantum devices, where controlling the directionality and energy states of photons is essential.
Next, we analyze the nonreciprocal absorption characteristics of the spatiotemporal quantum metasurface. The metasurface is designed to exhibit strong nonreciprocal behavior due to the space-time modulation, leading to direction-dependent absorption and transmission properties. Figures 4a and 4b illustrate the electric field distribution ($E_{y}$) and corresponding frequency spectrum for two different incident angles. Specifically, Figure 4a demonstrates the field distribution and frequency response when an electromagnetic wave is incident from the left at an angle of $55^{\circ}$. The results clearly show strong absorption of the incident wave by the metasurface, with minimal transmission, indicating effective photon blockade in this direction. The corresponding frequency spectrum confirms the absorption by showing the lack of transmitted energy at the incident frequency. Conversely, Figure 4b presents the scenario where the incident beam approaches from the right at an angle of $125^{\circ}$. In this case, the metasurface allows full transmission of the wave to the left side, effectively reflecting the wave back towards the source at an angle of $55^{\circ}$. The frequency spectrum corroborates this by displaying a significant transmitted signal at the original frequency, confirming the nonreciprocal behavior of the metasurface. Figures 4c and 4d provide the magnetic field distribution ($H_{z}$) under the same conditions. Figure 4c shows the absorption of the left-incident wave at $55^{\circ}$, while Figure 4d demonstrates the full transmission of the right-incident wave at $125^{\circ}$. The consistent behavior observed in both the electric and magnetic field distributions confirms the robustness of the nonreciprocal absorption mechanism induced by the spatiotemporal modulation in the superconducting quantum metasurface. These results demonstrate the capability of the designed metasurface to enforce nonreciprocal photon blockade, effectively controlling the directionality of absorption and transmission. The strong interaction between the incident photons and the spatiotemporally modulated nonlinear metasurface unit cells underpins the observed nonreciprocal behavior, which could be leveraged in various millikelvin-temperature quantum technology applications requiring directional control of photon propagation.
Figure 4: Nonreciprocal absorption in spatiotemporal quantum metasurfaces. (a) and (b) Field distribution (top) and frequency spectrum (bottom) for $E_{y}$, demonstrating (a) strong absorption of the incident beam from the left at $55^{\circ}$, and (b) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$. (c) and (d) Field distribution for $H_{z}$, demonstrating (c) strong absorption of the incident beam from the left at $55^{\circ}$, and (d) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.

### v4

@@
-Spatiotemporal Photon Blockade for Nonreciprocal Quantum Absorption
+Photon-Blockade Analogue Nonreciprocal Absorption in Spatiotemporal Metasurfaces
 Abstract
-Controlling the flow of photons is crucial for advancing quantum technologies.
-We introduce the concept of spatiotemporal photon blockade for nonreciprocal quantum absorption, utilizing space-time-periodic metasurfaces.
-Our study presents a methodology for experimentally realizing this effect, where photon frequency coherence with the metasurface’s space-time modulation enables one-way quantum absorption.
-In this system, forward-traveling photons are energetically modulated and absorbed within the slab, while backward-traveling photons are transmitted without interaction.
-Our analysis includes band structure, isofrequency diagrams, and nonreciprocal absorption results.
-These findings lay the groundwork for developing nonreciprocal quantum devices and enhancing photon management in milli-Kelvin temperature quantum systems.
-Photon blockade is a quantum optical phenomenon analogous to the Coulomb blockade, where the suppression of electrical transport occurs due to electronic repulsion, preventing multiple electrons from occupying the same region.
-In photon blockade, the absorption or transmission of a single photon inhibits the absorption or transmission of subsequent photons.
-This phenomenon occurs in systems where strong nonlinear interactions between photons effectively create a scenario in which only one photon can be absorbed or transmitted at a time, leading to quantized light behavior.
+Controlling the flow of electromagnetic energy is essential for advancing quantum technologies.
+We introduce a spatiotemporally modulated superconducting metasurface that exhibits photon-blockade-analogue nonreciprocal absorption.
+In this system, the frequency of incident radiation is matched to the modulation frequency of the metasurface, enabling one-way directional absorption.
+Forward-traveling waves undergo resonant coupling to higher-order Floquet harmonics and are absorbed within the slab, while backward-traveling waves transmit freely without interaction.
+This behavior arises from classical wave interference and harmonic conversion in a space-time periodic medium—a classical analogue of quantum photon blockade.
+We present a design based on a superconductor-semiconductor metasurface incorporating cascaded Josephson field-effect transistors (JoFETs) for millikelvin-temperature operation.
+Our analysis includes the system Hamiltonian, Floquet band structure, isofrequency diagrams, and full-wave simulations demonstrating strong nonreciprocal absorption.
+These findings establish a pathway toward compact, nonreciprocal superconducting devices for quantum information processing and microwave photonics.
+Photon blockade is a quantum optical phenomenon where the absorption or transmission of a single photon inhibits the absorption or transmission of subsequent photons.
 This effect, typically observed in systems such as atoms, quantum dots, or superconducting circuits coupled to resonant cavities, involves the system entering an excited state after the first photon interaction.
@@
 Observations of superconducting nonreciprocity through diamagnetic effects and engineered circuit asymmetries offer robust platforms for nonreciprocal quantum devices (?, ?, ?, ?).
-Traditional directional absorbers and electronic components, such as varactors, transistors, and diodes, face limitations and introduce noise at millikelvin temperatures typical of quantum technologies.
-To overcome these challenges, we propose a novel approach for nonreciprocal photon blockade through spatiotemporal modulation in quantum metasurfaces.
-By leveraging the unique properties of superconducting materials and dynamically modulating the metasurface structure, we achieve directional quantum absorption, where photons are selectively absorbed or transmitted based on their direction of incidence.
-This spatiotemporal photon blockade presents new opportunities for controlling quantum light-matter interactions and offers significant advancements for quantum information technologies and nonreciprocal photonic devices.
+Traditional directional absorbers and electronic components, such as varactors, transistors, and diodes, face limitations and introduce noise at millikelvin temperatures typical of superconducting quantum technologies.
+To overcome these challenges, we propose a spatiotemporally modulated superconducting metasurface that exhibits photon-blockade-analogue nonreciprocal absorption.
+By leveraging the unique properties of superconducting materials and dynamically modulating the metasurface structure, we achieve directional absorption where incident waves are selectively absorbed or transmitted based on their direction of incidence.
+This classical analogue of photon blockade, where forward-traveling waves undergo resonant harmonic conversion and absorption while backward-traveling waves transmit freely, offers new capabilities for controlling electromagnetic wave propagation in quantum-compatible platforms.
+The approach presents significant opportunities for advancing nonreciprocal devices and photon management in millikelvin-temperature quantum systems.
+Our approach involves nonlinear space-time-modulated metasurfaces incorporating gate-controlled Josephson field-effect transistors (JoFETs)  (?, ?, ?).
+These metasurfaces deliver highly efficient directional absorption while maintaining superior performance in superconducting quantum applications  (?).
+By enabling the miniaturization and integration of directional absorption components, our method preserves quantum state integrity and represents a significant leap forward in superconducting quantum technologies.
+The metasurface, featuring cascaded space-time-varying JoFETs, exploits the quantum mechanical phenomenon of supercurrent flow without resistance  (?, ?).
+This design harnesses the nonlinearity of JoFETs and their spatial and temporal modulation capabilities to advance directional absorption and wave reflection.
 Our approach achieves efficient directional absorption with simultaneous amplification, a rare combination in conventional devices.
 It addresses inefficiencies and spurious signal generation, paving the way for high-performance signal processing applications.
-Space-time refractive index modulation is conventionally realized through the modulation of electrical permittivity, as it is challenging to spatio-temporally modulate the magnetic permeability of materials  (?).
-In  (?, ?), it was shown for the first time that modulating both permeability and permittivity results in physical phenomena that cannot be achieved through conventional permittivity-modulated media, including zero reflection, zero band-gap, a large nonreciprocal response, and strong space-time harmonics.
-Furthermore, our metasurface demonstrates nonreciprocal behavior, with different absorption responses for incident waves from opposing directions, emphasizing the unidirectional properties imparted by the spatiotemporal modulation  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and time-varying systems (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?).
+Furthermore, our metasurface demonstrates nonreciprocal behavior, with different absorption responses for incident waves from opposing directions, emphasizing the unidirectional properties imparted by the spatiotemporal modulation (?, ?, ?, ?) and time-periodic devices (?, ?, ?, ?, ?, ?, ?, ?).
 Utilizing analytical techniques such as Bloch-Floquet solutions, we provide a comprehensive understanding of the wave scattering phenomena from our metasurface, validating its design and offering insights for further optimization.
-Unlike linear systems constrained by phase-matching and dispersion, our nonlinear spatiotemporal media achieve superior performance in a compact form, making them ideal for quantum technologies.
-1 Spatiotemporal Photon Blockade
-Consider photons with frequency $\omega_{0}$ incident on a slab with a space-time periodic current density $J(z,t)$, modulated at a temporal frequency $\omega_{\text{s}}$, where $\omega_{\text{s}}=\omega_{0}$.
-We demonstrate that the temporal coherence between the incident photons and the periodic modulation of the superconducting slab induces a nonreciprocal photon blockade.
+Unlike linear systems constrained by phase-matching and dispersion, our nonlinear superconducting spatiotemporal media achieve superior performance in a compact form, making them ideal for millikelvin-temperature superconducting quantum technologies.
+1 Photon-Blockade Analogue Mechanism for Nonreciprocal Absorption
+Figure 1: Nonreciprocal absorption in superconducting spatiotemporal metasurfaces, where the space-time modulation and incident wave share the same frequency ($\omega_{\text{s}}=\omega_{0}$).
+The system comprises space-time-modulated nonlinear unit cells coupled together.
+a) Incidence from the left port ($L_{\text{in}}$): Energy transition to higher states occur due to the strong interaction and one-way coherency with the nonlinear left-to-right traveling space-time modulation, resulting in zero transmission to the right port ($R_{\text{out}}=0$).
+b) Illustration of the physical mechanism for incidence from the right port ($R_{\text{in}}$): Free passage of wave to the left port ($L_{\text{out}}=R_{\text{in}}$) occurs due to the lack of transition to higher energy states and weak interaction with the opposite-direction traveling nonlinear space-time modulation, resulting in full transmission of waves.
+Consider waves with frequency $\omega_{0}$ incident on a superconducting slab with a space-time periodic current density $J(z,t)$, modulated at a temporal frequency $\omega_{\text{s}}$, where $\omega_{\text{s}}=\omega_{0}$.
+We demonstrate that the temporal coherence between the incident wave and the periodic modulation of the superconducting slab induces a nonreciprocal transmission.
 This phenomenon arises because the energy associated with the space-time modulation effectively inhibits the transmission of subsequent photons.
@@
 This results in a one-way quantum absorption process, facilitated by the coherent spatiotemporal modulation of the superconducting slab.
-Figure 1: Nonreciprocal photon blockade in superconducting spatiotemporal metasurfaces, where the space-time modulation and incident photons share the same frequency ($\omega_{\text{s}}=\omega_{0}$).
-The system comprises space-time-modulated nonlinear unit cells coupled together.
-a) Incidence from the left port ($L_{\text{in}}$): Photon blockade and energy transition to higher states occur due to the strong interaction and one-way coherency with the nonlinear left-to-right traveling space-time modulation, resulting in zero transmission to the right port ($R_{\text{out}}=0$).
-b) Illustration of the physical mechanism for incidence from the right port ($R_{\text{in}}$): Free passage of photons to the left port ($L_{\text{out}}=R_{\text{in}}$) occurs due to the lack of transition to higher energy states and weak interaction with the opposite-direction traveling nonlinear space-time modulation, resulting in full transmission of photons.
-Figure 1 illustrates the principle of spatiotemporal photon blockade enabling nonreciprocal quantum absorption by leveraging the interplay between the incident photons and the one-way coherent dynamic modulation.
+Figure 1 illustrates the principle of spatiotemporal nonreciprocal absorption by leveraging the interplay between the incident photons and the one-way coherent dynamic modulation.
 Figure 1a demonstrates the photon blockade effect when photons are incident from the left port ($L_{\text{in}}$).
@@
 In contrast, Fig.
-1b illustrates the scenario for photons incident from the right port ($R_{\text{in}}$).
+1b illustrates the scenario for wave incident from the right port ($R_{\text{in}}$).
 In this case, the interaction with the space-time modulation traveling in the opposite direction is weak.
-Consequently, there is minimal energy transition and the photons are transmitted freely to the left port ($L_{\text{out}}=R_{\text{in}}$).
-This nonreciprocal behavior, where the photon transmission and absorption are direction-dependent, underscores the potential of spatiotemporal metasurfaces for applications in quantum communication and sensing.
-Figure 2: One-way photon blockade and absorption functionality.
-Figure 2 illustrates the proposed one-way photon blockade and absorption mechanism facilitated by the coherently driven quantum spatiotemporal metasurface and its functionality.
-The metasurface operates as a nonreciprocal element, permitting photon transmission in one direction while blocking it in the opposite direction.
-To understand the nonreciprocal behavior of the spatiotemporal quantum metasurface, we first analyze its band structure using the dispersion relation.
+Consequently, there is minimal energy transition and the wave is transmitted freely to the left port ($L_{\text{out}}=R_{\text{in}}$).
+This nonreciprocal behavior, where the wave transmission and absorption are direction-dependent, underscores the potential of spatiotemporal metasurfaces for applications in quantum communication and sensing.
+Figure 2: Superconductor-semiconductor quantum spatiotemporal metasurface.
+(a) One-way absorption functionality.
+(b) Circuit model of the metasurface, consisting of an array of spatiotemporally-modulated gate-controlled Josephson field-effect transistors (JoFETs).
+Figure 2a illustrates the proposed one-way absorption mechanism facilitated by the quantum spatiotemporal metasurface and its functionality.
+The metasurface operates as a nonreciprocal element, permitting transmission in one direction while blocking it in the opposite direction.
+Figure 2b presents an equivalent circuit model of the metasurface, which comprises an array of space-time-periodic gate-controlled Josephson field-effect transistors (JoFETs).
+This circuit model is essential for understanding the underlying physical processes governing the nonreciprocal behavior.
+The current density across the spatiotemporal superconducting metasurface reads
+$J(z,t)=I_{0}\sin\left[\frac{2\pi\Phi(z,t)}{\Phi_{0}}\right],$ (1a)
+where $I_{0}$ is the maximum current, $\Phi(z,t)$ is the time- and space-dependent magnetic flux, and $\Phi_{0}$ is the magnetic flux quantum.
+Considering a voltage profile across the metasurface of the form $V(z,t)=V_{\text{dc}}+V_{\text{rf}}\sin[\kappa_{\text{s}}z-\omega_{\text{s}}t+\phi]$, where $V_{\text{dc}}$ is the direct current (DC) component, $V_{\text{rf}}$ is the radio frequency (RF) amplitude, $\kappa_{\text{s}}$ is the spatial wave number, $\omega_{\text{s}}$ is the modulation frequency, and $\phi$ is the phase.
+The corresponding space-time-periodic permeability of the JoFET array, which is critical for understanding the electromagnetic response of the metasurface, can be expressed as (Supplementary Text)
+$\mu_{\text{s}}(I,z,t)=\dfrac{l}{\mu_{0}A}\dfrac{\Phi_{0}}{2\pi I_{0}}\sec\left(\widetilde{\Phi}_{\text{dc}}+\widetilde{\Phi}_{\text{rf}}\sin[\kappa_{\text{s}}z-\omega_{\text{s}}t+\phi]\right),$ (1b)
+where $\widetilde{\Phi}_{\text{dc}}=\frac{2\pi\Phi_{\text{dc}}}{\Phi_{0}}$ and $\widetilde{\Phi}_{\text{rf}}=\frac{2\pi\Phi_{\text{rf}}}{\Phi_{0}}$ are the dimensionless forms of the DC and RF components of the magnetic flux, $l$ and $A$ are the length and the effective area associated with each JoFET, respectively, and $\mu_{0}$ is the permeability of free space.
+As the nonlinear permeability of JoFETs exhibits periodicity in both space and time, it can be represented by a Fourier series expansion, i.e.,
+$a(I,z,t)=\dfrac{1}{\mu_{\text{s}}(I,z,t)}=\sum_{m=-\infty}^{+\infty}a_{m}e^{-jm(\kappa_{\text{s}}z-\omega_{\text{s}}t+\phi)}.$ (2)
+We then determine the unknown coefficients in Eq. (2) using the Taylor series expansion.
+2 Theoretical Analysis
+2.1 Space-time-periodic superconducting metasurface
+As the nonlinear permeability of JoFETs in Eq. (1b) exhibits periodicity in both space and time, it can be represented by a Fourier series expansion, i.e.,
+$a(I,z,t)=\dfrac{1}{\mu_{\text{s}}(I,z,t)}=\sum_{m=-\infty}^{+\infty}a_{m}e^{-jm(\kappa_{\text{s}}z-\omega_{\text{s}}t+\phi)}.$ (3)
+We define $I_{\mu}=2\pi\mu_{0}I_{0}A/(\Phi_{0}l)$ and $\psi=\kappa_{\text{s}}z-\omega_{\text{s}}t+\phi$, and expand the cosine term
+$\begin{split}\cos\left(\widetilde{\Phi}_{\text{dc}}+\widetilde{\Phi}_{\text{rf}}\sin\psi\right)&=\cos\left(\widetilde{\Phi}_{\text{dc}}\right)\cos\left(\widetilde{\Phi}_{\text{rf}}\sin\psi\right)-\sin\left(\widetilde{\Phi}_{\text{dc}}\right)\sin\left(\widetilde{\Phi}_{\text{rf}}\sin\psi\right),\end{split}$ (4a)
+which can be further expanded using the Taylor’s series expansion, as
+$\begin{split}\cos&\left(\widetilde{\Phi}_{\text{rf}}\sin\psi\right)=\left(1-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{2}}}{4}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{4}}}{64}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{6}}}{2304}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{147456}\right)+\cos(2\psi)\left(\dfrac{{\widetilde{\Phi}_{\text{rf}}^{2}}}{4}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{4}}}{48}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{6}}}{1536}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{92160}\right)\\ &\qquad+\cos(4\psi)\left(\dfrac{{\widetilde{\Phi}_{\text{rf}}^{4}}}{192}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{6}}}{3840}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{184320}\right)+\cos(6\psi)\left(\dfrac{{\widetilde{\Phi}_{\text{rf}}^{6}}}{23040}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{645120}\right)+\cos(8\psi)\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{5160960},\end{split}$ (4b)
+$\begin{split}\sin\left(\widetilde{\Phi}_{\text{rf}}\sin\psi\right)=&\sin\psi\left(\widetilde{\Phi}_{\text{rf}}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{3}}}{8}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{5}}}{192}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{7}}}{9216}\right)+\sin(3\psi)\left(\dfrac{{\widetilde{\Phi}_{\text{rf}}^{3}}}{24}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{5}}}{384}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{7}}}{15360}\right)\\ &+\sin(5\psi)\left(\dfrac{{\widetilde{\Phi}_{\text{rf}}^{5}}}{1920}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{7}}}{46080}\right)-\sin(7\psi)\dfrac{{\widetilde{\Phi}_{\text{rf}}^{7}}}{322560},\end{split}$ (4c)
+where
+$\begin{split}&a_{0}=I_{\mu}\cos\left(\widetilde{\Phi}_{\text{dc}}\right)\left(1-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{2}}}{4}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{4}}}{64}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{6}}}{2304}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{147456}\right)\\ &a_{1}=-a_{-1}=\dfrac{I_{\mu}}{2j}\sin\left(\widetilde{\Phi}_{\text{dc}}\right)\left(\widetilde{\Phi}_{\text{rf}}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{3}}}{8}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{5}}}{192}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{7}}}{9216}\right)\\ &a_{2}=a_{-2}=\dfrac{I_{\mu}}{8}\cos\left(\widetilde{\Phi}_{\text{dc}}\right)\left({\widetilde{\Phi}_{\text{rf}}^{2}}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{4}}}{12}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{6}}}{384}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{23040}\right)\\ &a_{3}=-a_{-3}=\dfrac{I_{\mu}}{j48}\sin\left(\widetilde{\Phi}_{\text{dc}}\right)\left({\widetilde{\Phi}_{\text{rf}}^{3}}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{5}}}{16}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{7}}}{640}\right);a_{4}=a_{-4}=\dfrac{I_{\mu}}{384}\cos\left(\widetilde{\Phi}_{\text{dc}}\right)\left({\widetilde{\Phi}_{\text{rf}}^{4}}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{6}}}{20}+\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{960}\right)\\ &a_{5}=-a_{-5}=\dfrac{I_{\mu}}{j3840}\sin\left(\widetilde{\Phi}_{\text{dc}}\right)\left({\widetilde{\Phi}_{\text{rf}}^{5}}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{7}}}{24}\right);\quad a_{6}=a_{-6}=\dfrac{I_{\mu}}{46080}\cos\left(\widetilde{\Phi}_{\text{dc}}\right)\left({\widetilde{\Phi}_{\text{rf}}^{6}}-\dfrac{{\widetilde{\Phi}_{\text{rf}}^{8}}}{28}\right)\\ &a_{7}=-a_{-7}=\dfrac{I_{\mu}}{j645120}\sin\left(\widetilde{\Phi}_{\text{dc}}\right){\widetilde{\Phi}_{\text{rf}}^{7}};\quad a_{8}=a_{-8}=\dfrac{I_{\mu}}{10321920}\cos\left(\widetilde{\Phi}_{\text{dc}}\right){\widetilde{\Phi}_{\text{rf}}^{8}}.\end{split}$ (4d)
+2.2 Hamiltonian derivation
+For a single junction, the Josephson relations are given by
+$I_{j}=I_{c}(V_{g})\sin\phi_{j},\qquad V_{j}=\frac{\Phi_{0}}{2\pi}\frac{d\phi_{j}}{dt},$ (5)
+where $I_{c}(V_{g})$ is the gate-voltage-dependent critical current, $\phi_{j}$ is the superconducting phase difference across the $j$-th junction, and $\Phi_{0}=h/2e$ is the magnetic flux quantum.
+The energy stored in the junction comprises both the Josephson energy and the charging energy.
+The Josephson energy is
+$U_{j}=-E_{J}(V_{g})\cos\phi_{j},$ (6)
+with $E_{J}(V_{g})=\Phi_{0}I_{c}(V_{g})/(2\pi)$. The charging energy, associated with the junction capacitance $C_{j}$, takes the form
+$T_{j}=\frac{Q_{j}^{2}}{2C_{j}}=\frac{(2e)^{2}}{2C_{j}}n_{j}^{2}=4E_{C}n_{j}^{2},$ (7)
+where $Q_{j}=2en_{j}$ is the charge on the junction, $n_{j}$ is the number of Cooper pairs, and $E_{C}=e^{2}/(2C_{j})$ is the charging energy.
+The distinguishing feature of the proposed metasurface is the spatiotemporal modulation of the JoFET array via gate voltages.
+This modulation renders the Josephson energy both space- and time-dependent,
+$E_{J}(z,t)=E_{J,\text{dc}}+E_{J,\text{rf}}\sin(\kappa_{s}z-\omega_{s}t+\phi),$ (8a)
+and using the relation $E_{J}=(\Phi_{0}/2\pi)^{2}/L_{s}$, the modulated Josephson energy becomes
+$E_{J}(z,t)=\frac{\Phi_{0}I_{0}}{2\pi}\cos\left(\tilde{\Phi}_{\text{dc}}+\tilde{\Phi}_{\text{rf}}\sin(\kappa_{s}z-\omega_{s}t+\phi)\right).$ (8b)
+The discrete nature of the junction array can be extended to a continuum model, where the superconducting phase $\phi$ is promoted to a continuous field $\phi(z,t)$.
+The Hamiltonian density for the system is then
+$\mathcal{H}(z,t)=\frac{1}{2}C(z)\left(\frac{\Phi_{0}}{2\pi}\right)^{2}\left(\frac{\partial\phi}{\partial t}\right)^{2}+\frac{1}{2L(z,t)}\left(\frac{\Phi_{0}}{2\pi}\right)^{2}\left(\frac{\partial\phi}{\partial z}\right)^{2}-E_{J}(z,t)\cos\phi.$ (9)
+Quantization proceeds by elevating $\phi$ to an operator-valued field and expanding in plane-wave modes,
+$\hat{\phi}(z,t)=\sum_{k}\sqrt{\frac{\hbar}{2\omega_{k}C_{k}}}\left(\hat{a}_{k}e^{i(kz-\omega_{k}t)}+\hat{a}_{k}^{\dagger}e^{-i(kz-\omega_{k}t)}\right),$ (10)
+where $\hat{a}_{k}$ and $\hat{a}_{k}^{\dagger}$ are bosonic annihilation and creation operators satisfying $\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}]=\delta_{kk^{\prime}}$.
+The cosine term in the Hamiltonian introduces nonlinearity.
+For small phase fluctuations ($\phi\ll 1$), we expand
+$\cos\phi=1-\frac{\phi^{2}}{2}+\frac{\phi^{4}}{24}+\mathcal{O}(\phi^{6}).$ (11)
+The quadratic term yields the linear dynamics, while the quartic term gives rise to the nonlinear interactions essential for harmonic generation.
+Because the Josephson energy $E_{J}(z,t)$ is periodic in both space and time, it admits a Fourier series expansion,
+$E_{J}(z,t)=\sum_{m=-\infty}^{\infty}E_{J,m}e^{-im(\kappa_{s}z-\omega_{s}t+\phi)}.$ (12)
+Substituting the mode expansion and the Fourier representation of $E_{J}$ into the Hamiltonian, and retaining terms up to quartic order in the fields, yields
+$\mathcal{H}=\sum_{n,k}\hbar\omega_{k}\hat{a}_{n,k}^{\dagger}\hat{a}_{n,k}+\sum_{m}\hbar\omega_{m}\hat{b}_{m}^{\dagger}\hat{b}_{m}+\mathcal{H}_{\text{int}},$ (13)
+where $\hat{a}_{n,k}$ annihilates a photon in the $n$-th spatial harmonic with wavevector $k$, and $\hat{b}_{m}$ annihilates a modulation quantum in the $m$-th harmonic.
+The interaction Hamiltonian couples these modes via
+$\mathcal{H}_{\text{int}}=\sum_{m,n}g_{m,n}\left(\hat{a}_{n}^{\dagger}\hat{a}_{n+m}\hat{b}_{m}+\text{h.c.}\right).$ (14)
+In the regime where the modulation drive is strong (so that $\hat{b}_{m}$ can be treated classically as a large-amplitude coherent state, $\hat{b}_{m}\rightarrow\beta_{m}e^{-i\omega_{s}t}$), and considering only the coupling between the fundamental mode and the first modulation harmonic, the effective Hamiltonian simplifies to
+$\mathcal{H}_{\text{eff}}=\hbar\omega\hat{a}^{\dagger}\hat{a}+\sum_{m}\hbar\omega_{m}\hat{a}_{m}^{\dagger}\hat{a}_{m}+g\left(\hat{a}^{\dagger}\hat{a}_{m}e^{-i(\kappa_{s}z-\omega_{s}t+\phi)}+\hat{a}_{m}^{\dagger}\hat{a}e^{i(\kappa_{s}z-\omega_{s}t+\phi)}\right).$ (15)
+The connection to the classical wave equation used in our analysis follows from the Heisenberg equations of motion. For the operator $\hat{a}_{n}$, we have
+$i\hbar\frac{d\hat{a}_{n}}{dt}=[\hat{a}_{n},\mathcal{H}]=\hbar\omega_{n}\hat{a}_{n}+\sum_{m}g_{m,n}\hat{a}_{n+m}e^{-i(\kappa_{s}z-\omega_{s}t+\phi)}.$ (16)
+Taking the expectation value $\langle\hat{a}_{n}\rangle=\psi_{n}e^{-i\omega_{n}t}$ and assuming the fields are sufficiently coherent, we obtain
+$\omega_{n}\psi_{n}=\omega_{n}\psi_{n}+\sum_{m}g_{m,n}\psi_{n+m}e^{-i(\kappa_{s}z+\phi)},$ (17)
+which, after proper normalization, reduces to the matrix equation
+$\mathbf{A}]\cdot[\vec{\Psi}]=0.$ (18)
+2.3 Floquet space-time harmonics
+Given the spatiotemporal periodicity of the metasurface, the electric and magnetic fields within the slab can be expanded into Floquet space-time harmonics. The electric field can be expressed as
+$\begin{split}\mathbf{E}_{\text{s}}(x,z,t)&=\mathbf{\hat{y}}\sum_{n}\varPsi_{n}e^{-j\left[k_{x}x+\kappa_{n}z-\omega_{n}t\right]},\end{split}$ (19a)
+and the corresponding magnetic field is given by
+$\begin{split}&\mathbf{H}_{\text{s}}(x,z,t)=\frac{1}{\eta_{2}}\left[\mathbf{\hat{k}}_{\text{s}}\times\mathbf{E}_{\text{s}}(x,z,t)\right]=\dfrac{1}{\eta_{2}}\sum_{n}\left(-\mathbf{\hat{x}}\frac{\kappa_{n}}{k_{n}}+\mathbf{\hat{z}}\sin(\theta_{\text{i}})\right)\varPsi_{n}e^{-j\big[k_{x}x+\kappa_{n}z-\omega_{n}t\big]}.\end{split}$ (19b)
+Here, $\theta_{\text{i}}$ is the angle between the incident wave and the metasurface boundary, $\cos(\theta_{n})=\kappa_{n}/k_{n}$, $\mathbf{\hat{k}}_{\text{s}}=\mathbf{\hat{x}}\sin(\theta_{\text{i}})+\mathbf{\hat{z}}\kappa_{n}/k_{n}$, and $\kappa_{n}=\kappa_{0}+n\kappa_{\text{s}}$ is the $z$-component of the wavenumber for the $n$th space-time harmonic inside the metasurface.
+The source-free wave equation for the system is
+$\nabla^{2}\mathbf{H}_{\text{s}}(x,z,t)-\frac{1}{{{c^{2}}}}\frac{{{\partial^{2}}\left[\mu_{\text{s}}(z,t)\mathbf{H}_{\text{s}}(x,z,t)\right]}}{{\partial{t^{2}}}}=0.$ (20a)
+Substituting the magnetic field expression from Eq. (19b) into the wave equation (20a) yields
+$\begin{split}&\sum_{n}(k_{x}^{2}+\kappa_{n}^{2})\varPsi_{n}e^{-j\left[k_{x}x+\kappa_{n}z-\omega_{n}t\right]}-\dfrac{1}{c^{2}}\dfrac{\partial^{2}}{\partial{t^{2}}}\sum_{m,n}\widetilde{\mu}_{m}\varPsi_{m+n}e^{-j\left[k_{x}x+\kappa_{n}z-\omega_{n}t+\phi\right]}=0,\end{split}$ (20b)
+which may be cast as
+$\varPsi_{n}e^{j\phi}\left[\frac{k_{x}^{2}+\kappa_{n}^{2}}{k_{n}^{2}}\right]-\sum\limits_{m}\tilde{\mu}_{m}\varPsi_{m+n}=0,$ (20c)
+where $\tilde{\mu}_{m}=1/a_{m}$. Truncating to $2N+1$ terms yields
+$\mathbf{A}]\cdot[\vec{\Psi}]=0.$ (21a)
+where the matrix $\textbf{A}]$ is a square matrix of size $(2N+1)\times(2N+1)$, with elements defined as
+$\begin{split}A_{nn}&=\frac{k_{x}^{2}+\kappa_{n}^{2}}{k_{n}^{2}}-\tilde{\mu}_{0},\\ A_{nm}&=-\tilde{\mu}_{m+n},\quad\text{for }n\neq m.\end{split}$ (21b)
+The matrix $\textbf{A}]$ reads
+$\textbf{A}]=\begin{bmatrix}v_{-N}&\tilde{\mu}_{1}&\tilde{\mu}_{2}&\cdots&\tilde{\mu}_{M-2}&\tilde{\mu}_{M-1}&\tilde{\mu}_{M}\\ \tilde{\mu}_{-1}&v_{-N+1}&\tilde{\mu}_{1}&\cdots&\tilde{\mu}_{M-3}&\tilde{\mu}_{M-2}&\tilde{\mu}_{M-1}\\ \tilde{\mu}_{-2}&\tilde{\mu}_{-1}&v_{-N+2}&\cdots&\tilde{\mu}_{M-4}&\tilde{\mu}_{M-3}&\tilde{\mu}_{M-2}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \tilde{\mu}_{-M+2}&\tilde{\mu}_{-M+3}&\tilde{\mu}_{-M+4}&\cdots&v_{N-2}&\tilde{\mu}_{1}&\tilde{\mu}_{2}\\ \tilde{\mu}_{-M+1}&\tilde{\mu}_{-M+2}&\tilde{\mu}_{-M+3}&\cdots&\tilde{\mu}_{-1}&v_{N-1}&\tilde{\mu}_{1}\\ \tilde{\mu}_{-M}&\tilde{\mu}_{-M+1}&\tilde{\mu}_{-M+2}&\cdots&\tilde{\mu}_{-2}&\tilde{\mu}_{-1}&v_{N}\\ \end{bmatrix}$, (21c)
+where $v_{n}=\tilde{\mu}_{0}-(k_{x}^{2}+\kappa_{n}^{2})/k_{n}^{2}$.
+The vector of unknowns $\overrightarrow{\varPsi}]$ in Eq.
+(21a) is a $(2N+1)\times 1$ vector containing the $\varPsi_{n}$ values.
+For non-trivial solutions (i.e., $\overrightarrow{\varPsi}]\neq 0$), the matrix $\textbf{A}]$ must be singular, which implies that its determinant is zero, i.e.,
+$\det\left[\textbf{A}\right]=0.$ (22)
+The dispersion relation condition in Eq.
+(22) ensures that the system supports wave propagation, and provides the dispersion relation, which is typically expressed as $\omega(k)$ or $k(\omega)$.
+Matrix $\textbf{A}]$ contains material parameters, such as permittivity, permeability, and conductivity, which affect wave propagation.
+The matrix elements of $\textbf{A}]$ depend on the wavevector ($\kappa_{n}$) and frequency ($\omega_{n}$), establishing a connection between them, we find the conditions under which the system supports wave propagation.
+This relation describes how the frequency and wavevector are connected, namely, $\omega_{n}(\kappa_{n})$ or $\kappa_{n}(\omega_{n})$, and reveals how waves propagate through the medium, including information about allowed frequencies and wavevectors, group velocity and phase velocity, and band structures.
+Equation (22), along with the truncated $a_{n}$ coefficients (noting that, theoretically, an infinite number of elements exists), demonstrates the impact of nonlinearity, where energy from the primary harmonic modes ($A_{nn}$) is distributed across an infinite series of adjacent harmonic modes ($A_{nm}$, where $n\neq m$, $\infty<n,m<+\infty$).
+This contrasts with linear space-time-modulated media  (?, ?, ?), where energy from the main harmonic modes ($A_{nn}$) is distributed only to the immediately adjacent harmonic modes ($A_{n,n-1}$ and $A_{n,n+1}$).
+Nonlinear systems, however, possess the ability to redistribute energy across a broader spectrum of harmonic frequencies, allowing efficient transfer of energy from the incident wave to higher-frequency harmonic modes.
+This is because the nonlinear interactions can sustain the high-frequency modulation necessary for such transfers.
+We next consider incidence of transverse electromagnetic (TE) field
+$\mathbf{E}_{\text{I}}(x,z,t)=\mathbf{\hat{y}}E_{0}e^{j\omega_{0}t}e^{-j\left[k_{0}\sin(\theta_{\text{i}})x+k_{0}\cos(\theta_{\text{i}})z\right]}$ (23a)
+$\begin{split}\mathbf{H}_{\text{I}}(x,z,t)&=\frac{1}{\eta_{1}}\left[\mathbf{\hat{k}}_{\text{I}}\times\mathbf{E}_{\text{I}}(x,z,t)\right]\\ &=\frac{1}{\eta_{1}}\left[-\mathbf{\hat{x}}\cos(\theta_{\text{i}})+\mathbf{\hat{z}}\sin(\theta_{\text{i}})\right]E_{0}e^{j\omega_{0}t}e^{-j\left[k_{0}\sin(\theta_{\text{i}})x+k_{0}\cos(\theta_{\text{i}})z\right]}\end{split}$ (23b)
+where $\eta_{1}=\sqrt{\mu_{0}\mu_{\text{r}}/(\epsilon_{0}\epsilon_{r})}$. The reflected and transmitted electric fields outside of the slab may be defined as
+$\mathbf{E}_{\text{R}}(x,z,t)=\mathbf{\hat{y}}\sum\limits_{n=-\infty}^{\infty}R_{n}e^{-j\left[k_{0}\sin(\theta_{\text{i}})x-k_{0n}\cos(\theta^{\text{R}}_{n})z-\omega_{n}t\right]},$ (24a)
+$\begin{split}\mathbf{H}_{\text{R}}(x,z,t)&=\frac{1}{\eta_{1}}[\mathbf{\hat{k}}_{\text{R}}\times\mathbf{E}_{\text{R}}(x,z,t)]\\ &=\frac{1}{\eta_{1}}\left[\mathbf{\hat{x}}\cos(\theta^{\text{R}}_{n})+\mathbf{\hat{z}}\sin(\theta_{\text{i}})\right]R_{n}e^{-j\left[k_{0}\sin(\theta_{\text{i}})x-k_{0n}\cos(\theta^{\text{R}}_{n})z-\omega_{n}t\right]},\end{split}$ (24b)
+We then enforce the continuity of the tangential components of the electromagnetic fields at $z=0$ and $z=d$ to find the unknown field amplitudes $A_{0p}^{\pm}$, $R_{n}$ and $T_{n}$.
+The electric and magnetic fields continuity conditions between regions 1 and 2 at $z=0$, ${E_{1y}}(x,0,t)={E_{2y}}(x,0,t)$ and ${H_{1x}}(x,0,t)={H_{2x}}(x,0,t)$ read
+$\delta_{n0}E_{0}+R_{n}=\varPsi_{n},$ (25a)
+$\begin{split}&\cos(\theta_{\text{i}})\delta_{n0}E_{0}-\cos(\theta^{\text{R}}_{n})R_{n}=\dfrac{\eta_{1}}{\eta_{2}}\frac{\kappa_{n}}{k_{n}}\varPsi_{n},\end{split}$ (25b)
+where $k_{n}=(\omega_{0}+n\omega_{\text{s}})/v_{\text{b}}$.
+3 Results
+3.1 Unit-cell architecture
+To demonstrate the functionality of the proposed nonreciprocal superconducting quantum absorber metasurface, we design a metasurface with the thickness $d=0.6\lambda$, and the space-time-varying permeability given by $\mu_{\text{s}}(z,t)=\sec\left(0.6+0.75\sin[2\pi\times 4\times 10^{9}(1.67z/c-t)]\right)$.
+Figures 3a and 3b illustrate the experimental prototype design featuring a 2D array of JoFETs.
+Space-time modulation is achieved through a periodic signal traveling along the yellow lines and spatiotemporally modulate the gate voltage of JoFETs.
+Figure 3a illustrates a schematic of the space-time-modulated JoFET.
+The proposed spatiotemporal Josephson Field-Effect Transistor (JoFET) unit cell utilizes an Al-proximitized InAs quantum well, grown on a semi-insulating Fe counter-doped (100) InP wafer  (?), providing a voltage-controlled Josephson junction  (?, ?, ?).
+Figure 3: Experimental prototype design.
+(a) Schematic of the space-time-modulated superconductor-semiconductor Josephson field-effect transistors (JoFETs).
+(b) Metasurface architecture composing a 2D array of gated superconductor-semiconductor hybrid modulated JoFETs in space and time via the yellow-color modulation line.
+Given that the lattice parameter of the InAs active region is larger than that of the InP substrate, resulting in a lattice mismatch of approximately 3%, the formation of misfit and threading dislocations is inevitable.
+To mitigate this, we employ pseudomorphic growth using a step-graded buffer where the composition of the InxAl1-xAs buffer layer is gradually varied.
+The quantum well structure includes a 4 nm layer of InAs, followed by a 4 nm layer of In0.81Ga0.25As, and capped with a 10 nm layer of In0.81Ga0.25As.
+This design optimizes the interface quality while ensuring high mobility in the two-dimensional electron gas.
+[…]
+3.2 Dispersion analysis
+To understand the nonreciprocal behavior of the superconductor-semiconductor spatiotemporal quantum metasurface, we first analyze its band structure using the dispersion relation.
 The analysis reveals crucial insights into the energy transitions and nonreciprocity induced by the space-time modulation.
@@
 This diagram illustrates a strong nonreciprocity in the system.
-Specifically, the higher-order harmonics that travel in the forward $+z$ direction converge at the normalized wave vector $\kappa_{n}/\kappa_{\text{s}}=1$.
-This convergence indicates a resonant interaction between the incident photons and the space-time-modulated medium, leading to the absorption of energy from the fundamental harmonic.
+Specifically, the higher-order harmonics that travel in the forward +z+z direction converge at the normalized wave vector $\kappa_{n}/\kappa_{\text{s}}=1$.
+This convergence indicates a resonant interaction between the incident wave and the space-time-modulated medium, leading to the absorption of energy from the fundamental harmonic.
 The absorbed energy causes a transition from the fundamental frequency $\omega_{0}$ to higher energy states denoted as $\omega_{m}$.
-Figure 3: Band structure of the spatiotemporal quantum metasurface computed.
+Figure 4: Band structure of the superconductor-semiconductor spatiotemporal quantum metasurface computed using Eq.
+(22).
 (a) The $\omega-\kappa$ diagram illustrates strong nonreciprocity at $\omega_{0}/\omega_{\text{s}}=1$.
-Here, all higher-order forward ($+z$ direction) traveling space-time harmonics converge at $\kappa_{n}/\kappa_{\text{s}}=1$, resulting in the absorption of energy from the fundamental harmonic and transitioning the system from $\omega_{0}$ to higher frequency states, denoted as $\omega_{m}$.
-In contrast, the higher-order backward ($-z$ direction) traveling space-time harmonics are spatially separated, preventing any energy transition from $\omega_{0}$ to higher states.
+Here, all higher-order forward (+z+z direction) traveling space-time harmonics converge at $\kappa_{n}/\kappa_{\text{s}}=1$, resulting in the absorption of energy from the fundamental harmonic and transitioning the system from $\omega_{0}$ to higher frequency states, denoted as $\omega_{m}$.
+In contrast, the higher-order backward (z-z direction) traveling space-time harmonics are spatially separated, preventing any energy transition from $\omega_{0}$ to higher states.
 (b) The $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$ further demonstrates the nonreciprocity of the metasurface, where all forward higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$, effectively absorbing the energy of the fundamental harmonic.
-The group velocity vector $v_{\text{g},n}$ indicates that these harmonics propagate along the $+z$ direction, parallel to the metasurface boundary, rather than being transmitted outside the metasurface.
-In contrast, the higher-order harmonics that travel in the backward $-z$ direction are spatially separated in the $\omega-\kappa$ diagram.
-This separation prevents any significant interaction between the incident photons and the medium, thereby inhibiting the energy transition from $\omega_{0}$ to higher states.
+The group velocity vector $v_{\text{g},n}$ indicates that these harmonics propagate along the +z+z direction, parallel to the metasurface boundary, rather than being transmitted outside the metasurface.
+In contrast, the higher-order harmonics that travel in the backward z-z direction are spatially separated in the $\omega-\kappa$ diagram.
+This separation prevents any significant interaction between the incident wave and the medium, thereby inhibiting the energy transition from $\omega_{0}$ to higher states.
 As a result, the backward-propagating waves do not experience the same degree of energy absorption as the forward-propagating waves, which directly contributes to the observed nonreciprocal behavior of the metasurface.
@@
 This diagram further confirms the nonreciprocal nature of the metasurface, where all forward-propagating higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$.
-The clustering of these harmonics facilitates the absorption of the fundamental harmonic’s energy, while the group velocity vector $v_{\text{g},n}$ indicates that the energy is predominantly confined and directed along the $+z$ axis, parallel to the metasurface boundary.
+The clustering of these harmonics facilitates the absorption of the fundamental harmonic’s energy, while the group velocity vector $v_{\text{g},n}$ indicates that the energy is predominantly confined and directed along the +z+z axis, parallel to the metasurface boundary.
 This confinement prevents the transmission of energy to the exterior of the metasurface, leading to strong nonreciprocal absorption.
-The phenomena observed in the band structure is related to the concept of spatiotemporal photon blockade.
-In photon blockade, the presence of a single photon can prevent the passage of subsequent photons due to nonlinear interactions, effectively ”blocking” further photon transmission.
-In our system, this blockade effect is realized through the spatiotemporal modulation of the metasurface.
+The phenomena observed in the band structure is related to the concept of spatiotemporal photon blockade, where the presence of a single photon can prevent the passage of subsequent photons due to nonlinear interactions, effectively ”blocking” further photon transmission.
+In our system, this effect is realized through the spatiotemporal modulation of the metasurface.
 The space-time modulation, with a frequency $\omega_{\text{s}}$ matched to the incident photon frequency $\omega_{0}$, induces a strong nonlinear interaction for forward-propagating waves.
 As a result, when photons from the left are incident on the metasurface, they are resonantly coupled to the modulation, leading to energy absorption and transition to higher frequency states (spatial harmonics) $\omega_{m}$.
-This absorption effectively prevents the transmission of these photons through the metasurface, analogous to a photon blockade.
+This absorption effectively prevents the transmission through the metasurface.
 The nonreciprocity arises because the same modulation does not facilitate similar energy transitions for backward-propagating waves, allowing them to pass through the metasurface without significant energy loss.
-Thus, the metasurface behaves as a nonreciprocal photon blockade device, where the space-time modulation selectively blocks photons traveling in one direction by shifting their energy to higher states, while allowing photons traveling in the opposite direction to pass freely.
+Thus, the metasurface behaves as a nonreciprocal device, where the space-time modulation selectively blocks traveling in one direction by shifting their energy to higher states, while allowing wave traveling in the opposite direction to pass freely.
 This behavior is crucial for potential applications in quantum information processing and nonreciprocal quantum devices, where controlling the directionality and energy states of photons is essential.
-Next, we analyze the nonreciprocal absorption characteristics of the spatiotemporal quantum metasurface.
+3.3 Transmission and fields distributions
+Figure 5: Nonreciprocal absorption in spatiotemporal superconductor-semiconductor quantum metasurfaces.
+(a) and (b) Field distribution (top) and frequency spectrum (bottom) for $E_{y}$, demonstrating (a) strong absorption of the incident beam from the left at $55^{\circ}$, and (b) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.
+(c) and (d) Field distribution for $H_{z}$, demonstrating (c) strong absorption of the incident beam from the left at $55^{\circ}$, and (d) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.
+Next, we analyze the nonreciprocal absorption characteristics of the spatiotemporal superconductor-semiconductor quantum metasurface.
 The metasurface is designed to exhibit strong nonreciprocal behavior due to the space-time modulation, leading to direction-dependent absorption and transmission properties.
-Figures 4a and 4b illustrate the electric field distribution ($E_{y}$) and corresponding frequency spectrum for two different incident angles.
+Figures 5a and 5b illustrate the electric field distribution ($E_{y}$) and corresponding frequency spectrum for two different incident angles.
 Specifically, Figure 5a demonstrates the field distribution and frequency response when an electromagnetic wave is incident from the left at an angle of $55^{\circ}$.
-The results clearly show strong absorption of the incident wave by the metasurface, with minimal transmission, indicating effective photon blockade in this direction.
+The results clearly show strong absorption of the incident wave by the metasurface, with minimal transmission.
 The corresponding frequency spectrum confirms the absorption by showing the lack of transmitted energy at the incident frequency.
@@
 The frequency spectrum corroborates this by displaying a significant transmitted signal at the original frequency, confirming the nonreciprocal behavior of the metasurface.
-Figures 4c and 4d provide the magnetic field distribution ($H_{z}$) under the same conditions.
+Figures 5c and 5d provide the magnetic field distribution ($H_{z}$) under the same conditions.
 Figure 5c shows the absorption of the left-incident wave at $55^{\circ}$, while Figure 5d demonstrates the full transmission of the right-incident wave at $125^{\circ}$.
 The consistent behavior observed in both the electric and magnetic field distributions confirms the robustness of the nonreciprocal absorption mechanism induced by the spatiotemporal modulation in the superconducting quantum metasurface.
-These results demonstrate the capability of the designed metasurface to enforce nonreciprocal photon blockade, effectively controlling the directionality of absorption and transmission.
-The strong interaction between the incident photons and the spatiotemporally modulated nonlinear metasurface unit cells underpins the observed nonreciprocal behavior, which could be leveraged in various millikelvin-temperature quantum technology applications requiring directional control of photon propagation.
-Figure 4: Nonreciprocal absorption in spatiotemporal quantum metasurfaces.
-(a) and (b) Field distribution (top) and frequency spectrum (bottom) for $E_{y}$, demonstrating (a) strong absorption of the incident beam from the left at $55^{\circ}$, and (b) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.
-(c) and (d) Field distribution for $H_{z}$, demonstrating (c) strong absorption of the incident beam from the left at $55^{\circ}$, and (d) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.
+These results demonstrate the capability of the designed metasurface to enforce nonreciprocity, effectively controlling the directionality of absorption and transmission.
+The strong interaction between the incident photons and the spatiotemporally modulated nonlinear metasurface unit cells underpins the observed nonreciprocal behavior, which could be leveraged in various millikelvin-temperature quantum technology applications requiring directional control of wave propagation.